# On Bits and Bandits: Quantifying the Regret-Information Trade-off

### v3

@@ 1 Introduction
In interactive decision-making problems, an agent repeatedly interacts with a task by sequentially choosing decisions from a decision space. Subsequently, the agent receives feedback that usually includes a reward and, optionally, other types of information. For example, the feedback includes only the reward in multi-armed bandit (MAB) problems [Lattimore and Szepesvári, 2020]. In partial monitoring, however, it only includes a signal, which the agent then utilizes to indirectly infer the reward [Cesa-Bianchi et al., 2006]. The goal of the agent in such tasks is to minimize the gap between the accumulated reward and the reward of the best decision in hindsight, also known as regret.
In these tasks, agents can accumulate information from multiple sources, that can be categorized into two main types. One is information acquired through direct interactions and receiving feedback from the task. For example, in reinforcement learning (RL) such information includes visited states, accumulated rewards, and performed actions. The other type is information provided by external sources, such as human advice [Najar and Chetouani, 2021], text description generated by large language models [Du et al., 2023, Hu and Sadigh, 2023], and more.
Resource allocation is an example of a task that combines both types of information. In this task, our goal is to route a process from a mobile device to a server while minimizing energy consumption and execution time [Zuo and Joe-Wong, 2021]. In this task, we can accumulate information directly by routing processes and receiving feedback. However, we also have relevant information that can be obtained from other sources, such as server specifications, load information, and more.
Consider two agents playing on the same online task. Both agents have access to the same observations, but the first agent has already played multiple rounds in the task. Thus, he already suffered regret and gained an advantage over the second agent. Before he starts interacting with the task, the second agent can use external knowledge that will allow him to reach the same performance level as the first agent. In this paper, we focus on how much information the second agent needs to query so he will "catch up" with the first agent without suffering regret. In other words, we focus on the following question:
@@ 1.1 Contributions
We introduce a novel information-theoretic method to obtain regret lower bounds for a wide variety of interactive decision-making problems. This method uses Fano’s inequality [Cover, 1999] and the construction method introduced by Yang and Barron [1999] to lower bound the worst-case Bayesian regret. We then use this method to derive lower bounds for MAB, contextual bandit, and RL tasks.
@@
We consider an interactive decision-making framework that describes general online tasks with contextual information in a Bayesian setting. We examine the regret of policies with a mutual information constraint and obtain Bayesian regret lower bounds that depend on the amount of information an agent accumulates. Additionally, we show upper bounds for Thompson sampling [Thompson, 1933] in an MAB environment, which depends on the information the agent collects. We also present lower bounds for scenarios where the entropy of the prior is constrained, for a specific family of algorithms.
@@ 2 Setting and Preliminaries
We introduce a new setting called interactive decision-making in Bayesian environments with contextual information, which is illustrated in Figure 1. This setting is a Bayesian adaptation of the frequentist interactive decision-making setting introduced by Foster et al. [2021]. We denote the set of Borel probability measures over a locally compact space by $\Delta(\cdot)$. Let $\Pi$ be a compact decision space, $\mathcal{M}$ be a compact model space and $\mathcal{C}$ be a compact context space. Every model $M\in\mathcal{M}$ is a mapping from the decision and context space to an observation space $\mathcal{O}$, i.e., $M:\Pi\times\mathcal{C}\to\mathcal{O}$. Every model $M$ has a reward function, $R_{M}:\Pi\times\mathcal{C}\to\Delta([0,1])$ that maps every decision-context pair to a reward distribution over $[0,1]$. A task in this setting is defined by the decision space $\Pi$, the model class $\mathcal{M}$, the context space $\mathcal{C}$, the reward functions $R_{M}$, and the prior $P\in\Delta(\mathcal{M}\times\mathcal{C})$. We denote by $\mu_{M}(\pi,C)=\mathbb{E}_{x\sim R_{M}(\pi,C)}[x]$ the mean reward of decision $\pi$ under model $M$ and context $C$. With a slight abuse of notation, we also denote by $\mu_{M}(\phi,C)=\mathbb{E}_{\pi\sim\phi}[\mu_{M}(\pi,C)]$, the mean reward of a policy $\phi\in\Delta(\Pi)$ under model $M$ and context $C$.
In a $T$-round game, a random model $M$ is sampled before the game starts according to the prior $P$. Then a context $C_{t}$ is sampled every round, independently from previous rounds, according to the marginal prior probability $P(\cdot\mid M)$. Before each round starts, the context $C_{t}$ is revealed to the agent, who determines his policy $p_{t}:\mathcal{C}\to\Delta(\Pi)$ according to the context and the observed history. Then, a decision $\pi_{t}\sim p_{t}$ is sampled and an observation $O_{t}\sim M(\pi_{t},C_{t})$ is revealed to the agent. A reward, $r_{t}\sim R_{M}(\pi_{t},C_{t})$ is also generated. We denote the history up to time $t$ as $\mathcal{H}_{t}=\{C_{i},\pi_{i},O_{i}\}_{i=1}^{t}$. We measure the performance with respect to the best context-dependent decision in hindsight. This performance metric is called the Bayesian regret and is defined by
$\mathcal{BR}_{P}(T;\{p_{t}\}_{t=1}^{T})=\sup_{\pi^{*}:\mathcal{C}\to\Pi}\mathbb{E}\left[\sum_{t=1}^{T}\left(\mu_{M}(\pi^{*}({C_{t}}),C_{t})-r_{t}\right)\right]$ (1)
@@
The contextual MAB problem is defined by the rewards of each arm, which are, say, Bernoulli random variables. For simplicity, we assume a finite set of decisions, which are called arms in this setting, $\Pi=\{1,2,\dots,K\}$. Furthermore, we assume that the context space is finite as well, $\mathcal{C}=\{1,2,\dots,C\}$. The model class is then given by =$\mathcal{M}=\{f|f:\Pi\times\mathcal{C}\to[0,1]\}$, so every model maps $\left(\pi,C\right)$ to $\mathrm{Bern}(f(\pi,C))$. At each round $t$, a context is sampled and revealed. A decision $\pi_{t}$ is selected and the observation is the incurred reward, $r_{t}=O_{t}\sim\mathrm{Bern}(f_{t}(\pi_{t},C_{t}))$.
@@
Consider a normed space $A$ with metric $d$ induced by the norm, and let $\epsilon>0$. We denote by $\mathcal{N}(d,A,\epsilon)$ and $\mathcal{M}(d,A,\epsilon)$ the $\epsilon$-covering and $\epsilon$-packing numbers of $A$ under the norm $d$, respectively. We define an $\epsilon$-local set of $A$ as $B(b,\epsilon)=\{a\in A:d(b,a)\leq\epsilon\}$, for every $b\in A$ and $\epsilon>0$. The local packing number, denoted by $\mathcal{M}^{\mathrm{loc}}\left(d,A,\epsilon\right)$, is the largest $(\epsilon/2)$-packing set of any set $B(b,\epsilon)$, i.e.,
@@
We denote by $H(X)=\mathbb{E}\left[-\log Q(X)\right]$ the Shannon entropy of a random variable $X\sim Q$, and with a slight abuse of notation, we also denote $H(Q)=H(X)$. Given two jointly distributed random variables, $(X,Y)\sim Q$, we denote their marginal distributions by $Q(x)$ and $Q(y)$. We also define the mutual information of two jointly distributed random variables $(X,Y)\sim Q$ as $I(X;Y)=\mathbb{E}\left[\log\left(\frac{Q(X,Y)}{Q(X)Q(Y)}\right)\right]$. The KL-divergence is defined as $D\left(P,Q\right)=\mathbb{E}_{X\sim P}\left[\log\frac{P(X)}{Q(X)}\right]$ [Cover, 1999]. Unless stated otherwise, logarithms are assumed to be in base-2, denoted by $\log$. We also denote $x\gtrsim y$ if there is some global constant $c>0$ such that $x\geq cy$. We define by $\mathbb{B}_{p}^{d}=\{x\in\mathbb{R}^{d}:\left\|x\right\|_{p}\leq 1\}$ the unit sphere in $\mathbb{R}^{d}$ under the $L_{p}$ norm.
@@
MAB $\Omega(\sqrt{KT})$ Yes [Audibert and Bubeck, 2009]
Tabular RL $\Omega(\sqrt{HSKT})$ Yes [Zhang et al., 2021]
@@
Lipschitz bandit $\Omega\left(T^{\frac{d+1}{d+2}}\right)$ Yes [Kleinberg et al., 2019]
@@ 3 Regret Lower Bounds Using Fano’s Inequality
We now introduce a novel approach for obtaining worst-case Bayesian regret lower bounds, which can be applied to a wide variety of interactive decision-making tasks. This method can be viewed as an extension of Fano’s inequality for non-exact recovery [Scarlett and Cevher, 2019] to online problems. We define worst-case Bayesian regret in the following manner,
@@
Theorem 3.1 (Fano’s inequality, Theorem 2.10 of Cover [1999]).
@@
While Theorem 3.1 is used to lower-bound the error in hypothesis testing problems, we now show how it can also be applied to regret minimization. We assume access to some set $\Phi=\{\phi_{1},\dots,\phi_{K}\}$ of (possibly stochastic) policies. We also assume that a metric $\rho$ exists such that $\mathbb{E}[\left|\mu_{M}(\psi,C)-\mu_{M}(\phi,C)\right|]\geq\rho(\phi,\psi)$ for all $\psi,\phi\in\Delta(\Pi)$. The following proposition applies Theorem 3.1 to regret minimization, by reducing it to a hypothesis testing over the finite set $\Phi$.
@@
The proof of Proposition 3.2 is provided in Appendix C.1. Proposition 3.2 allows us to lower-bound the Bayesian regret using a specific set of decisions. We suggest a method for selecting the decision set $\Phi$, which provides a general approach for obtaining regret lower bounds. We separate our analysis into finite, parametric with infinite decisions, and non-parametric decision spaces. This separation is done to make the lower bound in Equation 4 tighter. We observe that the lower bound is tighter for larger values of $K$. Thus, we should select the largest value of $K$ that satisfies the requirement $\rho(\phi_{i},\phi_{j})\geq\epsilon$ for a given $\epsilon>0$. For finite decision spaces, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision simplex, $\Delta(\Pi)$, according to $\rho$, under $P$. So, we have $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)$. Similarly, for parametric decision spaces with infinite decisions, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision set, $\Pi$, i.e., $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Pi,\epsilon)$. For non-parametric decision spaces, we use an $\epsilon$-global packing set of our decision space $\Pi$, which means that $K=\mathcal{M}(\rho,\Pi,\epsilon)$. After selecting $\Phi$, we need to handle $I(V;\mathcal{H}_{T})$. To do this, we make the following assumption about the policies.
@@
There exist constants $\bar{A},\epsilon_{0}>0$ such that for all policies $p_{1},p_{2}$ such that $\rho(p_{1},p_{2})\leq\epsilon_{0}$, $D\left(p_{1},p_{2}\right)\leq 2\bar{A}\rho(p_{1},p_{2})^{2}$.
This assumption is also made by Yang and Barron [1999]. The following theorem upper bounds $I(V;\mathcal{H}_{T})$.
Theorem 3.3 (Yang and Barron [1999]).
@@
The proof of Theorem 3.4 is provided in Appendix C.2. To obtain regret lower bounds, all we need are the values of the packing or local packing, and covering numbers of the decision space. Then, we simply need to select a value of $\epsilon$ such that the right-hand side of Equations 7, 8 or 9 is a positive constant. In Appendix C.3 we demonstrate this, by deriving regret lower bounds for several known online problems, which are summarized in Table 1. Moreover, this method can be used to obtain lower bounds for the frequentist regret, due to the mini-max theorem, which states that the worst-case Bayesian regret is equal to the mini-max frequentist regret [Lattimore and Szepesvári, 2019].
4 Information Theoretic Bayesian Regret Lower Bounds
@@ 4.1 Mutual Information Constraint
We now derive Bayesian regret bounds under a constraint on the total amount of information that the agent accumulates. Intuitively, in information-theoretic terms, mutual information quantifies the amount of information one random variable contains about another one [Cover, 1999]. This means that mutual information can be utilized to measure the information that is aggregated from various sources using a common measure of bits. We are interested in lower bounding the regret, under a constraint on the amount of information the agent accumulates. Then, we ask the following question: What is the lower bound of the worst-case Bayesian regret, when the agent accumulates less than $R$ bits of information? We will use the method described in Section 3 to derive regret lower bounds under this constraint. We note that in general $I(\pi^{*};\mathcal{H}_{T})=D\left(P^{\pi^{*}\mid\mathcal{H}_{T}},P^{\pi^{*}}\right)$. Thus, for general agents, $D\left(p_{t},P^{\pi^{*}}\right)$ is the information accumulated up to round $t$, in bits. The following proposition answers this question, for problems with a finite decision space.
@@
Let $R\leq\log K$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, such that $D\left(p_{t},P^{\pi^{*}}\right)\leq R$ for all $t$. Then,
@@
Assume an $K$-MAB problem with no constraints. Let $\tilde{O}_{t},\tilde{\mathcal{H}}_{T}$ be constrained observations and history such that $I(\pi^{*};\tilde{\mathcal{H}_{T}})=R$. The regret of a policy that utilizes Thompson sampling with these observations is upper bounded by $O\left(\log K\sqrt{\frac{KT}{R}}\right)$. Additionally, if the MAB is linear the regret can be upper bounded by $O\left(\log K\sqrt{\frac{dT}{R}}\right)$.

### v4

@@ 1 Introduction
In interactive decision-making problems, an agent repeatedly interacts with a task by sequentially choosing decisions from a decision space. Subsequently, the agent receives feedback that usually includes a reward and, optionally, other types of information. For example, the feedback includes only the reward in multi-armed bandit (MAB) problems (Lattimore & Szepesvári, 2020). In partial monitoring, however, it only includes a signal, which the agent then utilizes to indirectly infer the reward (Cesa-Bianchi et al., 2006). The goal of the agent in such tasks is to minimize the gap between the accumulated reward and the reward of the best decision in hindsight, also known as regret.
In these tasks, agents can accumulate information from multiple sources, that can be categorized into two main types. One is information acquired through direct interactions and receiving feedback from the task. For example, in reinforcement learning (RL) such information includes visited states, accumulated rewards, and performed actions. The other type is information provided by external sources, such as human advice (Najar & Chetouani, 2021), text description generated by large language models (Du et al., 2023; Hu & Sadigh, 2023), and more.
In some tasks both information from direct interactions and external information are present. One example is RL with LLM (Du et al., 2023), where an agent interacts with an online environment and is provided with advice from an LLM. The agent can interact with the environment as well as ask the LLM for advice.
Consider two agents playing on the same online task. Both agents have access to the same observations, but the first agent has already played multiple rounds in the task. Thus, he already suffered regret and gained an advantage over the second agent. Before he starts interacting with the task, the second agent can use external knowledge that will allow him to reach the same performance level as the first agent. In this paper, we focus on how much information the second agent needs to query so he will ”catch up” with the first agent without suffering regret. In other words, we focus on the following question:
@@ 1.1 Contributions
We consider an interactive decision-making framework that describes general online tasks with contextual information in a Bayesian setting. We then introduce a novel information-theoretic method to obtain regret lower bounds for decision-making problems in this setting. This method uses Fano’s inequality (Cover, 1999) and the construction method introduced by Yang & Barron (1999) to lower bound the worst-case Bayesian regret. In Theorem 3.4 we introduce the general lower bound, and Table 1 shows its application on MAB, contextual bandit, and RL tasks.
@@
The total amount of information an agent gathers can be quantified to bits by the mutual information functional. Proposition 4.1 presents Bayesian regret lower bounds that depend on the amount of information an agent accumulates. Additionally, Proposition 4.3 upper bounds the regret for Thompson sampling (Thompson, 1933) in an MAB environment, which depends on the information the agent collects. We also present lower bounds for scenarios where the entropy of the prior is constrained in Proposition 4.5.
@@ 2 Setting and Preliminaries
We introduce a new setting called interactive decision-making in Bayesian environments with contextual information, which is illustrated in Figure 1. This setting is a Bayesian adaptation of the frequentist interactive decision-making setting introduced by Foster et al. (2021). We denote the set of Borel probability measures over a locally compact space by $\Delta(\cdot)$. Let $\Pi$ be a compact decision space, $\mathcal{M}$ be a compact model space and $\mathcal{C}$ be a compact context space. Every model $M\in\mathcal{M}$ is a mapping from the decision and context space to an observation space $\mathcal{O}$, i.e., $M:\Pi\times\mathcal{C}\to\mathcal{O}$. Every model $M$ has a reward function, $R_{M}:\Pi\times\mathcal{C}\to\Delta([0,1])$ that maps every decision-context pair to a reward distribution over $[0,1]$. A task in this setting is defined by the decision space $\Pi$, the model class $\mathcal{M}$, the context space $\mathcal{C}$, the reward functions $R_{M}$, and the prior $P\in\Delta(\mathcal{M}\times\mathcal{C})$. We denote by $\mu_{M}(\pi,C)=\mathbb{E}_{x\sim R_{M}(\pi,C)}[x]$ the mean reward of decision $\pi$ under model $M$ and context $C$. With a slight abuse of notation, we also denote by $\mu_{M}(\phi,C)=\mathbb{E}_{\pi\sim\phi}[\mu_{M}(\pi,C)]$, the mean reward of a stochastic decision $\phi\in\Delta(\Pi)$ under model $M$ and context $C$.
In a $T$-round game, a random model $M$ is sampled before the game starts according to the prior $P$. Then a context $C_{t}$ is sampled every round, independently from previous rounds, according to the marginal prior probability $P(\cdot\mid M)$. Before each round starts, the context $C_{t}$ is revealed to the agent, who determines his stochastic decision $p_{t}:\mathcal{C}\to\Delta(\Pi)$ according to the context and the observed history. Then, a decision $\pi_{t}\sim p_{t}$ is sampled and an observation $O_{t}\sim M(\pi_{t},C_{t})$ is revealed to the agent. A reward, $r_{t}\sim R_{M}(\pi_{t},C_{t})$ is also generated. We denote the history up to time $t$ as $\mathcal{H}_{t}=\{C_{i},\pi_{i},O_{i}\}_{i=1}^{t}$. We denote $\pi^{*}_{M,C}=\arg\max_{\pi\in\Pi}\mu_{M}(\pi,C)$. We measure the performance with respect to the best context-dependent decision in hindsight. This performance metric is called the Bayesian regret and is defined by
$\mathcal{BR}_{P}(T;\{p_{t}\}_{t=1}^{T})=\mathbb{E}\left[\sum_{t=1}^{T}\left(\mu_{M}({\pi^{*}_{M,C_{t}}},C_{t})-r_{t}\right)\right]$ (1)
@@
The contextual MAB problem is defined by the rewards of each arm, which are, say, Bernoulli random variables. For simplicity, we assume a finite set of decisions, which are called arms in this setting, $\Pi=\{1,2,\dots,K\}$. Furthermore, we assume that the context space is finite as well, $\mathcal{C}=\{1,2,\dots,C\}$. The model class is then given by =$\mathcal{M}=\{f|f:\Pi\times\mathcal{C}\to[0,1]\}$, so every model maps $\left(\pi,C\right)$ to $\mathrm{Bern}(f(\pi,C))$. At each round $t$, a context is sampled and revealed. A decision $\pi_{t}$ is selected and the observation is the incurred reward, $r_{t}=O_{t}\sim\mathrm{Bern}(f_{t}(\pi_{t},C_{t}))$. We note that the Bernoulli rewards were selected to maintain a simple example, and that our setting also covers general reward distributions.
@@
Consider a normed space $A$ with metric $d$ induced by the norm, and let $\epsilon>0$. We denote by $\mathcal{N}(d,A,\epsilon)$ and $\mathcal{M}(d,A,\epsilon)$ the $\epsilon$-covering and $\epsilon$-packing numbers of $A$ under the norm $d$, respectively. We define an $\epsilon$-ball of $b\in A$ as $B(b,\epsilon)=\{a\in A:d(b,a)\leq\epsilon\}$, for every $b\in A$ and $\epsilon>0$. The local packing number, denoted by $\mathcal{M}^{\mathrm{loc}}\left(d,A,\epsilon\right)$, is the largest $(\epsilon/2)$-packing set of any set $B(b,\epsilon)$, i.e.,
@@
We denote by $H(X)=\mathbb{E}\left[-\log Q(X)\right]$ the Shannon entropy of a random variable $X\sim Q$, and with a slight abuse of notation, we also denote $H(Q)=H(X)$. Given two jointly distributed random variables, $(X,Y)\sim Q$, we denote their marginal distributions by $Q(x)$ and $Q(y)$. We also define the mutual information of two jointly distributed random variables $(X,Y)\sim Q$ as $I(X;Y)=\mathbb{E}\left[\log\left(\frac{Q(X,Y)}{Q(X)Q(Y)}\right)\right]$. The KL-divergence is defined as $D_{\mathrm{KL}}\left(P\parallel Q\right)=\mathbb{E}_{X\sim P}\left[\log\frac{P(X)}{Q(X)}\right]$ (Cover, 1999). Unless stated otherwise, logarithms are assumed to be in base-2, denoted by $\log$. We also denote $x\gtrsim y$ if there is some global constant $c>0$ such that $x\geq cy$. We define by $\mathbb{B}_{p}^{d}=\{x\in\mathbb{R}^{d}:\left\|x\right\|_{p}\leq 1\}$ the unit sphere in $\mathbb{R}^{d}$ under the $L_{p}$ norm.
@@
MAB $\Omega(\sqrt{KT})$ Yes (Audibert & Bubeck, 2009)
Tabular RL $\Omega(\sqrt{HSKT})$ Yes (Zhang et al., 2021)
@@
Lipschitz bandit $\Omega\left(T^{\frac{d+1}{d+2}}\right)$ Yes (Kleinberg et al., 2019)
@@ 3 Regret Lower Bounds Using Fano’s Inequality
We now introduce a novel approach for obtaining worst-case Bayesian regret lower bounds, which can be applied to a wide variety of interactive decision-making tasks. This method can be viewed as an extension of Fano’s inequality for non-exact recovery (Scarlett & Cevher, 2019) to online problems. We define worst-case Bayesian regret in the following manner,
@@
Theorem 3.1 (Fano’s inequality, Theorem 2.10 of Cover (1999)).
@@
While Theorem 3.1 is used to lower-bound the error in hypothesis testing problems, we now show how it can also be applied to regret minimization. We assume access to some set $\Phi=\{\phi_{1},\dots,\phi_{K}\}$ of (possibly stochastic) decisions. We also assume that a metric $\rho$ exists such that $\mathbb{E}[\left|\mu_{M}(\psi,C)-\mu_{M}(\phi,C)\right|]\geq\rho(\phi,\psi)$ for all $\psi,\phi\in\Delta(\Pi)$. The following proposition applies Theorem 3.1 to regret minimization, by reducing it to a hypothesis testing over the finite set $\Phi$.
@@
The proof of Proposition 3.2 is provided in Appendix C.1. Proposition 3.2 allows us to lower-bound the Bayesian regret using a specific set of decisions. We suggest a method for selecting the decision set $\Phi$, which provides a general approach for obtaining regret lower bounds. We separate our analysis into finite, parametric with infinite decisions, and non-parametric decision spaces. This separation is done to make the lower bound in Equation 4 tighter. We observe that the lower bound is tighter for larger values of $K$. Thus, we should select the largest value of $K$ that satisfies the requirement $\rho(\phi_{i},\phi_{j})\geq\epsilon$ for a given $\epsilon>0$. For finite decision spaces, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision simplex, $\Delta(\Pi)$, according to $\rho$, under $P$. So, we have $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)$. Similarly, for parametric decision spaces with infinite decisions, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision set, $\Pi$, i.e., $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Pi,\epsilon)$. For non-parametric decision spaces, we use an $\epsilon$-global packing set of our decision space $\Pi$, which means that $K=\mathcal{M}(\rho,\Pi,\epsilon)$. After selecting $\Phi$, we need to handle $I(V;\mathcal{H}_{T})$. To do this, we make the following assumption about the decisions.
@@
There exist constants $\bar{A},\epsilon_{0}>0$ such that for all stochastic decisions $p_{1},p_{2}$ such that $\rho(p_{1},p_{2})\leq\epsilon_{0}$, $D_{\mathrm{KL}}\left(p_{1}\parallel p_{2}\right)\leq 2\bar{A}\rho(p_{1},p_{2})^{2}$.
This assumption means that for all stochastic decisions in an $\epsilon_{0}$-ball around $p_{1}$, the KL divergence can be upper bounded using $\rho$, and is commonly made, for example, by Yang & Barron (1999). The following theorem upper bounds $I(V;\mathcal{H}_{T})$.
Theorem 3.3 (Yang & Barron (1999)).
@@
The proof of Theorem 3.4 is provided in Appendix C.2. To obtain regret lower bounds, all we need are the values of the packing or local packing, and covering numbers of the decision space. Then, we simply need to select a value of $\epsilon$ such that the right-hand side of Equations 7, 8 or 9 is a positive constant. In Appendix C.3 we demonstrate this, by deriving regret lower bounds for several known online problems, which are summarized in Table 1. Moreover, this method can be used to obtain lower bounds for the frequentist regret, due to the mini-max theorem, which states that the worst-case Bayesian regret is equal to the mini-max frequentist regret (Lattimore & Szepesvári, 2019).
4 Information Theoretic Bayesian Regret Upper and Lower Bounds
@@ 4.1 Mutual Information Constraint
We now derive Bayesian regret bounds under a constraint on the total amount of information that the agent accumulates. Intuitively, in information-theoretic terms, mutual information quantifies the amount of information one random variable contains about another one (Cover, 1999). This means that mutual information can be utilized to measure the information that is aggregated from various sources using a common measure of bits. We are interested in lower bounding the regret, under a constraint on the amount of information the agent accumulates. Then, we ask the following question: What is the lower bound of the worst-case Bayesian regret, when the agent accumulates less than $R$ bits of information? We will use the method described in Section 3 to derive regret lower bounds under this constraint. We note that in general $I(\pi^{*};\mathcal{H}_{T})=D_{\mathrm{KL}}\left(P^{\pi^{*}\mid\mathcal{H}_{T}}\parallel P^{\pi^{*}}\right)$. Thus, for general agents, $D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)$ is the information accumulated up to round $t$, in bits. The following proposition answers this question, for problems with a finite decision space.
@@
Let $R\leq\log K$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, such that $D_{\mathrm{KL}}\left(p_{t}\parallel P^{\pi^{*}}\right)\leq R$ for all $t$. Then,
@@
Assume an $K$-MAB problem with no constraints. Let $\tilde{O}_{t},\tilde{\mathcal{H}}_{T}$ be constrained observations and history such that $I(\pi^{*};\tilde{\mathcal{H}_{T}})=R$. The regret of Thompson sampling with these observations is upper bounded by $O\left(\log K\sqrt{\frac{KT}{R}}\right)$. Additionally, if the MAB is linear the regret can be upper bounded by $O\left(\log K\sqrt{\frac{dT}{R}}\right)$.